1. five times the sum of twice a number and 5 is the same as the sum of 9 times the number and 5.
the unknown number is?

Found 2 solutions by Edwin McCravy, ikleyn:
Answer by Edwin McCravy(20055)

five times the sum of twice a number and 5 is the same as the sum of 9 times the number and 5.
Change "a number" and "the number" to N.

five times the sum of twice N and 5 is the same as the sum of 9 times N and 5.
Change "twice N" to 2N

five times the sum of 2N and 5 is the same as the sum of 9 times N and 5.
Change "9 times N" to 9N

five times the sum of 2N and 5 is the same as the sum of 9N and 5.
Change "the sum of 2N and 5" to (2N+5) 

five times (2N+5) is the same as the sum of 9N and 5.
Change "the sum of 9N and 5" to 9N+5

five times (2N+5) is the same as 9N+5.

5(2N+5) is the same as 9N+5.
Change "is the same as" to an equal sign =.

5(2N+5) = 9N+5.
Now you can finish.

Edwin

Answer by ikleyn(52781)

.
The equation is

5*(2x + 5) = 9x + 5.

Solve it:

10x + 25 = 9x + 5,

10x - 9x = 5-25 = -20.

x = -20.

Check:  The left side is  5*(2x + 5) = 5*(2*(-20) + 5) = 5*(-40 + 5) = 5*(-35) = -175.

             The right side is  9x + 5 = 9*(-20) + 5 = -180 + 5 = -175.